OpenStudy (swamijatin2):

two years ago ,father was three times as old as his son and two year hence,twice his age will be equal to five times that of his son.find the peresent ages.

OpenStudy (swamijatin2):

please help

Directrix (directrix):

I am working on some equations now.

OpenStudy (swamijatin2):

ok

Directrix (directrix):

I started with this but something is not right. Maybe you can see the error in the set-up of the equations. Let F be the father's age today Let S be the son's age today Father's age 2 years ago: F - 2 Son's age 2 years ago: S - 2 So, F - 2 = 3*(S - 2) and ...

Directrix (directrix):

Father's age 2 years from today: F + 2 Son's age 2 years from today: S + 2 F + 2 = 5*(S + 2)

Directrix (directrix):

F - 2 = 3*(S - 2) F + 2 = 5*(S + 2) ============ Solving these simultaneously yields F = - 22 and S = -6 which cannot be correct. I need to re-read the question.

Directrix (directrix):

>> two year hence I assume that is two years from today and not two years from two years ago because that would be today. :)

OpenStudy (swamijatin2):

wait i m telling u the answer

OpenStudy (swamijatin2):

28 yrs,14yrs

Directrix (directrix):

I am trying to re-do my equations. Something is wrong with my work.

OpenStudy (swamijatin2):

in my school note book two years ago age is taken as father- 3x+2 son- x+2 but why this happen i dont know

mathslover (mathslover):

two years ago ,father was three times as old as his son and two year hence,twice his age will be equal to five times that of his son.find the peresent ages. Let fathers' present age is x and son's age be y 2 years ago, fathers' age was: x-2 = 3y 2 years after, father's age will be : x + 2 => 2(x+2) = 5y

mathslover (mathslover):

hmm.. something is wrong :D

Directrix (directrix):

>>in my school note book two years ago age is taken as father- 3x+2 son- x+2 father-> 3x+2 ==> This is father's age today, then son->x+2 ==> This is son's age today, then

mathslover (mathslover):

oops.. got it! 2(x+2) = 5(y+2) x-2 = 3(y-2) => x -2 = 3y - 6 => x = 3y -4 2(3y-4+2) = 5(y+2) 2(3y-2) = 5(y+2) 6y - 4 = 5y + 10 y = 14

OpenStudy (swamijatin2):

but why we take +2

OpenStudy (swamijatin2):

in question there is ago

mathslover (mathslover):

Directrix will help u out. :)

OpenStudy (swamijatin2):

vo to pta ni kha gya